OpenStudy (anonymous):

consider the linear equation -3x-4y=-1 find the y intercept and x intercept for the equation

OpenStudy (anonymous):

x intercept-when y=0 y intercept-when x=0 plug 0 in for y -3x-4(0)=-1 x=? plug 0 in for x -3(0)-4y=-1 y=?

OpenStudy (anonymous):

I keep getting a fraction

OpenStudy (anonymous):

can u tell me what answer you get for this problem please

OpenStudy (anonymous):

thats fine...it could be a whole number or a fraction. what answers did you get?

OpenStudy (anonymous):

I got 0/1 for the y intercept

OpenStudy (anonymous):

Ok, lets work through the x intercept. -3x-4(0)=-1 -3x=-1 divide both sides by -3 x=-1/-3 x=1/3 so the x intercept is (1/3, 0)

OpenStudy (anonymous):

now try to solve for y in this equation -3(0)-4y=-1

OpenStudy (anonymous):

y=0,1/4

OpenStudy (anonymous):

yes you get y=1/4 so you're right that makes the y intercept (0, 1/4)

OpenStudy (anonymous):

THANK U

OpenStudy (anonymous):

You're welcome